Russell invests $8100 in two different accounts. The first account paid 4 %, the second account paid 10 % in interest. At the end of the first year he had earned $504 in interest. How much was in each account?
$  at 4 %
$  at 10 %

Answer by macston(5194)

.
F=amount at 4%; T=amount at 10%
.
F+T=$8100
T=$8100-F
.
0.04F+0.10T=$504
0.04F+0.10($8100-F)=$504
0.04F+$810-0.10F=$504
-0.06F=-$306
F=$5100 ANSWER 1: $5100 was in the 4% account.
.
T=$8100-F=$8100-$5100=$3000 ANSWER 2: $3000 was in the 10% account.
.
CHECK:
0.04F+0.10T=$504
0.04($5100)+0.10($3000)=$504
$204+$300=$504
$504=$504